The velocity of an object is given as a function of time by v=8t+6t^2-16t^3. v is in m/s and time is in sec. Its average velocity over the time interval from t=0 to t=3?
Answer #1

Given equation v 8t 6t2 -16t at timet 0, velocity of the object is (0 s)-8 8(0) 6 (0) -16(0) at time t3, velocity of the obje
Answer #2


we can get the average velocity from   the eqn (V1+V2)/2

here V1 is at t = 0   and is

V1 = 8*0 + 6 *0^2   -16 * 0^3

V1 = 0

similarly V2 (at t   = 3 ) is

V2 = (8*3)   +(6* 3*3) - (16* 3*3*3)

V2 = -354 m//s

so avg velocoity    V = 0+ (-354)/3

V =-118 m/s----<<<<<<<<<<<<<<<< is the answer